Mathematics 241 – Multivariable Calculus Quiz 1 – September 10, 2010

Your Name:

A) (10) Describe in words the set described by S = {(x, y, z) ∈ R^{3} | x^{2} + y^{2} + z^{2} + 2z = 0} (Hint: complete the square in z).

B) (10) Give a set notation description for the line in R^{3} through the point p = (1, 2, −1) in the direction of the vector v = (4, 1, 7).